-5x -2y =10
y =(4x-11)

Answer by mananth(16946)

We solve by substitution method
-4 x + 1 y = -11 .............1
1 y = -11 + 4 x
/ 1 =
y = -11 + 4 x
-5 x + -2 y = 10 .....................2
Substitute y in (2)
-5 x + -2 ( -11 + 4 x ) = 10
-5 x + 22 + -8 x = 10
-5 x -8 x = -12
-13 x = -12
/ -13
x= 12/13
Plug the value of x in (1)
-4 x + 1 y = -11
-4 * 12/13 + 1 y = -11
-48/13 + 1 y = -11
1 y = -95/13
y= -95/13
m.ananth@hotmail.ca